# Characterization of flip process rules with the same trajectories

### v2

@@ Abstract.
Garbe, Hladký, Šileikis and Skerman [Ann. Inst. Henri Poincaré Probab. Stat., 60 (2024), pp. 2878–2922] recently introduced a general class of random graph processes called flip processes and proved that the typical evolution of these discrete-time random graph processes correspond to certain continuous-time deterministic graphon trajectories. We obtain a complete characterization of the equivalence classes of flip process rules with the same graphon trajectories. As an application, we characterize the flip process rules which are unique in their equivalence classes. These include several natural families of rules such as the complementing rules, the component completion rules, the extremist rules, and the clique removal rules.
@@ 1. Introduction
Graphs are mathematical structures which underpin the study of numerous real-life settings, many of which evolve in time according to a stochastic rule. This leads us to random graph processes, the study of which was initiated by Erdős and Rényi in their seminal paper [5] from 1960. In the decades that followed, a variety of random graph processes were introduced and studied extensively. One example is the triangle removal process, which was introduced by Bollobás and Erdős [4] in 1990. The process begins with the initial graph $G_{0}=K_{n}$, and in each step $\ell$ we obtain the graph $G_{\ell}$ by deleting the edges of a uniformly random triangle in $G_{\ell-1}$. The process terminates when $G_{\ell}$ is triangle-free. There is a long line of research on random graph processes with triangle-free final graphs, which has led to progressive breakthroughs on lower bounds for the Ramsey numbers $R(3,t)$ in [8, 2, 3, 6].
@@
Generalising the idea of local replacements, Garbe, Hladký, Šileikis and Skerman [7] recently introduced a general class of random graph processes called flip processes. Before we describe these processes, we first introduce some notation and terminology. Write $\mathbb{N}$ for the set of positive integers and $\mathbb{N}_{0}$ for the set $\mathbb{N}\cup\{0\}$. For $a\in\mathbb{N}_{0}$ write $[a]$ for the set $\{1,\dots,a\}$ and $[a]_{0}$ for $[a]\cup\{0\}$. All our graphs are simple and undirected. Let $k\in\mathbb{N}$ and let $\mathcal{H}_{k}$ be the set of all labelled graphs on the vertex set $[k]$. A rule is a matrix $\mathcal{R}=(\mathcal{R}_{F,H})_{F,H\in\mathcal{H}_{k}}\in[0,1]^{\mathcal{H}_{k}\times\mathcal{H}_{k}}$ such that for each $F\in\mathcal{H}_{k}$ we have $\sum_{H\in\mathcal{H}_{k}}\mathcal{R}_{F,H}=1$. We call $k$ the order of the rule $\mathcal{R}$. A flip process with rule $\mathcal{R}$ is a random graph process $(G_{\ell})_{\ell\in\mathbb{N}_{0}}$ where we start with an initial graph $G_{0}$ on the vertex set $[n]$ with $n\geq k$ and in step $\ell\in\mathbb{N}$ we obtain the graph $G_{\ell}$ from $G_{\ell-1}$ as follows. First, we sample an ordered tuple $\mathbf{v}=(v_{i})_{i\in[k]}$ of distinct vertices in $[n]$ at random. Next, we define the drawn graph $F\in\mathcal{H}_{k}$ by $ij\in E(F)$ if and only if $v_{i}v_{j}\in E(G_{\ell-1})$ and generate the replacement graph $H\in\mathcal{H}_{k}$ according to the probability distribution $(\mathcal{R}_{F,H})_{H\in\mathcal{H}_{k}}$. Finally, we replace $G_{\ell-1}[\mathbf{v}]$ with $H$ to obtain $G_{\ell}$ from $G_{\ell-1}$.
Flip processes were studied in [7] through the lens of dense graph limits [9]. Let $(\Omega,\pi)$ be an atomless probability space with an implicit separable sigma-algebra. Write $\mathcal{W}\subseteq L^{\infty}(\Omega^{2})$ for the set of kernels, that is, bounded symmetric measurable functions. Write $\mathcal{W}_{0}$ for the set of graphons, that is, the elements of $\mathcal{W}$ whose range is a subset of $[0,1]$. An important metric on $\mathcal{W}$ is the cut norm distance $d_{\square}$ given by $d_{\square}(U,W)=\sup_{S,T\subseteq\Omega}\left|\int_{S\times T}(U-W)\,\mathrm{d}\pi^{2}\right|$. Let $G$ be a graph and $(\Omega_{i})_{i\in V(G)}$ be a partition of $\Omega$ into parts of equal measure. We write $W_{G}$ for the graphon which is equal to $\mathbb{I}_{\left\{ij\in E(G)\right\}}$ on $\Omega_{i}\times\Omega_{j}$ and call it a graphon representation of $G$.
@@
We say that two rules $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ of orders $k_{1}$ and $k_{2}$ respectively have the same flip process distributions if for all finite graphs $G$ on $n\geq\max\{k_{1},k_{2}\}$ vertices the flip process started at $G$ with rule $\mathcal{R}_{1}$ has the same distribution as the flip process started at $G$ with rule $\mathcal{R}_{2}$. Note that two rules with the same trajectories could have different flip process distributions. A rule $\mathcal{R}$ is ignorant if the replacement graph distribution $(\mathcal{R}_{F,H})_{H\in\mathcal{H}_{k}}$ is independent of the drawn graph $F\in\mathcal{H}_{k}$. It was observed in [1, Section 4] that two ignorant rules $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ of the same order and with the same expected replacement edge count $d_{\mathcal{R}}=\sum_{H\in\mathcal{H}_{k}}\mathcal{R}_{F,H}e(H)$ have the same trajectories, that is, $\Phi_{\mathcal{R}_{1}}=\Phi_{\mathcal{R}_{2}}$. More concretely, consider ignorant rules $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ of order $4$ where for $\mathcal{R}_{1}$ the replacement graph is either complete or empty with probability $\frac{1}{2}$ each and for $\mathcal{R}_{2}$ the replacement graph is a copy of $K_{1,3}$. Clearly, they have different replacement behaviour. On the other hand, both rules are ignorant rules of the same order with the same expected replacement edge count $d_{\mathcal{R}_{1}}=d_{\mathcal{R}_{2}}=3$, so they have the same trajectories. Indeed, the lens of dense graph limits captures the macroscopic profile of flip processes but is insensitive to lower order fluctuations. This leads us to the following question.
@@
Here $T_{F^{a,b}}W(x,y)$ is invariant under the relabelling of vertices of $F^{a,b}$ (including $a$ and $b$), so we may focus on equivalence classes of rooted graphs $F^{a,b}$ related by relabelling; write $\mathcal{J}_{k}$ for the set of such equivalence classes for $\mathcal{G}_{k}$. For each $J\in\mathcal{J}_{k}$ let $T_{J}W(x,y)$ be the unique value of $T_{F^{a,b}}W(x,y)$ for all $F^{a,b}\in J$ and set $q_{J,\mathcal{R}}=\sum_{F^{a,b}\in J}q_{F^{a,b},\mathcal{R}}$. Hence, we have
@@
Theorem 1.3 characterizes equivalence classes of rules of the same order with the same trajectories. Let us now demonstrate how it may be applied to determine whether two rules $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ of different orders have the same trajectories. Say the orders of $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are $k_{1}$ and $k_{2}$ respectively; by symmetry we may assume $k_{1}<k_{2}$. Let $\mathcal{R}^{*}_{1}$ be the rule of order $k_{2}$ where the replacement graph $H$ is obtained from the drawn graph $F$ by replacing the subgraph of $F$ induced on $[k_{1}]$ according to rule $\mathcal{R}_{1}$ and leaving all other pairs unchanged. Since $\mathcal{R}^{*}_{1}$ acts exclusively on pairs on $[k_{1}]$ and on those pairs it acts in accordance with $\mathcal{R}_{1}$, it follows that $\mathcal{R}^{*}_{1}$ and $\mathcal{R}_{1}$ have the same trajectories. Let $F^{a,b}$ be a rooted graph with $F\in\mathcal{H}_{k_{2}}$ and distinct $a,b\in[k_{2}]$. If $a,b\in[k_{1}]$, then we have $q_{F^{a,b},\mathcal{R}^{*}_{1}}=q_{G^{a,b},\mathcal{R}_{1}}$ with $G$ being the subgraph of $F$ induced on $[k_{1}]$. Otherwise, we have $q_{F^{a,b},\mathcal{R}^{*}_{1}}=0$. Now we apply Theorem 1.3 to determine whether $\Phi_{\mathcal{R}^{*}_{1}}=\Phi_{\mathcal{R}_{2}}$, which is equivalent to $\Phi_{\mathcal{R}_{1}}=\Phi_{\mathcal{R}_{2}}$. All in all, this gives us a complete characterization of equivalence classes of flip process rules with the same trajectories.
Many combinatorially interesting rules turn out to be symmetric deterministic rules. Write $S_{k}$ for the symmetric group on $[k]$. We have the following natural action of $S_{k}$ on $\mathcal{H}_{k}$. Given $\sigma\in S_{k}$ and $F\in\mathcal{H}_{k}$ we define $\sigma\cdot F$ to be the graph with vertex set $[k]$ and edge set $\{\sigma(i)\sigma(j):ij\in E(F)\}$. A rule $\mathcal{R}$ of order $k$ is symmetric if for all $F,H\in\mathcal{H}_{k}$ and $\sigma\in S_{k}$ we have $\mathcal{R}_{\sigma\cdot F,\sigma\cdot H}=\mathcal{R}_{F,H}$ and deterministic if for each $F\in\mathcal{H}_{k}$ there is a unique $G\in\mathcal{H}_{k}$ such that $\mathcal{R}_{F,H}=\mathbb{I}_{\left\{H=G\right\}}$. Indeed, naturally arising rules are likely to act symmetrically on isomorphic graphs. For example, naturally arising rules of order $3$ are likely to treat the cherry graph $K_{1,2}$ symmetrically, no matter whether the middle vertex is labelled $1$, $2$ or $3$. The class of symmetric deterministic rules contains many natural families of rules which were highlighted in [7] and studied in [1]. These include the complementing rules, the component completion rules, the extremist rules, and the clique removal rules.
@@ 2. Main Theorem
    There is an open interval $\mathcal{D}_{\mathcal{R},W}\subseteq\mathbb{R}$ containing 00 and a trajectory $\Phi_{\mathcal{R}}^{\cdot}W\colon\mathcal{D}_{\mathcal{R},W}\to\mathcal{W}$ starting at $W$ such that any other trajectory starting at $W$ is a restriction of $\Phi_{\mathcal{R}}^{\cdot}W$ to a subinterval of $\mathcal{D}_{\mathcal{R},W}$.
@@
    If $W\in\mathcal{W}_{0}$ is a graphon, then the set $L_{\mathcal{R},W}:=\{t\in\mathcal{D}_{\mathcal{R},W}:\Phi_{\mathcal{R}}^{t}W\in\mathcal{W}_{0}\}$ is a closed interval containing $[0,\infty)$.
We introduce some further notation, terminology and concepts. Let $\mathbf{F}$ and $\mathbf{H}$ be rooted graphs. A function $\phi\colon V(\mathbf{F})\to V(\mathbf{H})$ is relation-preserving if for all $u,v\in V(\mathbf{F})$ we have $uv\in E(\mathbf{F})$ if and only if $\phi(u)\phi(v)\in E(\mathbf{H})$, root-respecting if for all $v\in V(\mathbf{F})$ we have $\phi(v)\in R(\mathbf{H})$ if and only if $v\in R(\mathbf{F})$, and root-order-preserving if for all $u,v\in R(\mathbf{F})$ we have $\phi(u)\leq_{R(\mathbf{H})}\phi(v)$ if and only if $u\leq_{R(\mathbf{F})}v$. A function $\phi\colon V(\mathbf{F})\to V(\mathbf{H})$ is nearly induced copy if it is relation-preserving and root-respecting; it is induced copy if it is additionally root-order-preserving. We say that $\mathbf{F}$ and $\mathbf{H}$ are isomorphic if there is an isomorphism from $\mathbf{F}$ to $\mathbf{H}$, that is, a bijective induced copy function $\phi\colon V(\mathbf{F})\to V(\mathbf{H})$; write $\mathbf{F}\cong\mathbf{H}$. Finally, let $\operatorname{Aut}(\mathbf{F})$ denote the set of automorphisms of $\mathbf{F}$, that is, isomorphisms from $\mathbf{F}$ to itself.
Let $k\in\mathbb{N}$. Write $\mathcal{J}_{k}$ for the set of isomorphism classes of the collection $\mathcal{G}_{k}$. For $J\in\mathcal{J}_{k}$ and a rule $\mathcal{R}$ of order $k$, set
@@
    For all $J\in\mathcal{J}_{k}$ we have $q_{J,\mathcal{R}_{1}}=q_{J,\mathcal{R}_{2}}$
@@
    $\mathfrak{V}_{\mathcal{R}_{1}}=\mathfrak{V}_{\mathcal{R}_{2}}$ on $\mathcal{W}_{0}$
@@
    $\mathfrak{V}_{\mathcal{R}_{1}}=\mathfrak{V}_{\mathcal{R}_{2}}$ on $\mathcal{W}$
@@
In this section we prove Theorem 2.3. We briefly describe our proof strategy here. Our proof scheme is (S1) $\Rightarrow$ (S3) $\Rightarrow$ (S4) $\Rightarrow$ (S2) $\Rightarrow$ (S1). We note that rooted induced densities are invariant under rooted graph isomorphism, so for $J\in\mathcal{J}_{k}$ we may set $T_{J}$ to be the unique operator equal to $T_{F^{a,b}}$ for all $F^{a,b}\in J$ and the first implication follows. The second implication follows from the definition and uniqueness of trajectories (Definition 2.1 and Theorem 2.2), while the third implication follows from the smoothness of the velocity operator (Lemmas 3.3 and 3.4). Indeed, the final implication (S2) $\Rightarrow$ (S1) is the most difficult step. The goal here is to show that the collection =$\mathcal{I}=\{T_{J}:J\in\mathcal{J}_{k}\}$ of operators is linearly independent.
We apply a couple of key insights to achieve this goal. Firstly, we focus on parametrized graphon representations of graphs because evaluating induced densities at these graphons nicely discretizes them into sums of monomials indexed over induced copy functions between rooted graphs. Secondly, we focus on twinfree versions of rooted graphs because Lemma 3.1 tells us that the aforementioned functions are necessarily injective when the graphs involved are twinfree; in particular, this induces a partial order on the set of twinfree versions of rooted graphs. Combining these two insights leads us to Lemma 3.2, which provides an explicit quantification of induced densities at parametrized graphon representations of graphs with a unique monomial assigned to each relevant rooted graph. Finally, the desired linear independence follows from the linear independence of monomials in polynomial functions (see Lemma 3.5).

### v3

@@ Abstract.
Garbe, Hladký, Šileikis and Skerman [Ann. Inst. Henri Poincaré Probab. Stat., 60 (2024), pp. 2878–2922] recently introduced a general class of random graph processes called flip processes and proved that the typical evolution of these discrete-time random graph processes corresponds to certain continuous-time deterministic graphon trajectories. We obtain a complete characterization of the equivalence classes of flip process rules with the same graphon trajectories. As an application, we characterize the flip process rules which are unique in their equivalence classes. These include several natural families of rules such as the complementing rules, the component completion rules, the extremist rules, and the clique removal rules.
@@ 1. Introduction
Graphs are mathematical structures which underpin the study of numerous real-life settings, many of which evolve in time according to a stochastic rule. This leads us to random graph processes, the study of which was initiated by Erdős and Rényi in their seminal paper [5] from 1960. In the decades that followed, a variety of random graph processes were introduced and studied extensively. One example is the triangle removal process, which was introduced by Bollobás and Erdős [4] in 1990. The process begins with the initial graph $G_{0}=K_{n}$, and in each step $\ell$ we obtain the graph $G_{\ell}$ by deleting the edges of a uniformly random triangle in $G_{\ell-1}$. The process terminates when $G_{\ell}$ is triangle-free. There is a long line of research on random graph processes resulting in triangle-free graphs, which has led to a series of breakthroughs on lower bounds for the Ramsey numbers $R(3,t)$ in [8, 2, 3, 6].
@@
Generalising the idea of local replacements, Garbe, Hladký, Šileikis and Skerman [7] recently introduced a general class of random graph processes called flip processes. Before we describe these processes, we first introduce some notation and terminology. Write $\mathbb{N}$ for the set of positive integers and $\mathbb{N}_{0}$ for the set $\mathbb{N}\cup\{0\}$. For $a\in\mathbb{N}_{0}$ write $[a]$ for the set $\{1,\dots,a\}$ and $a]_{0}$ for $a]\cup\{0\}$. All our graphs are simple and undirected. Let $k\in\mathbb{N}$ and let $\mathcal{H}_{k}$ be the set of all labelled graphs on the vertex set $[k]$. A rule is a matrix $\mathcal{R}=(\mathcal{R}_{F,H})_{F,H\in\mathcal{H}_{k}}\in[0,1]^{\mathcal{H}_{k}\times\mathcal{H}_{k}}$ such that for each $F\in\mathcal{H}_{k}$ we have $\sum_{H\in\mathcal{H}_{k}}\mathcal{R}_{F,H}=1$. We call $k$ the order of the rule $\mathcal{R}$. A flip process with rule $\mathcal{R}$ is a random graph process $(G_{\ell})_{\ell\in\mathbb{N}_{0}}$ where we start with an initial graph $G_{0}$ on the vertex set $[n]$ with $n\geq k$ and in step $\ell\in\mathbb{N}$ we obtain the graph $G_{\ell}$ from $G_{\ell-1}$ as follows. First, we sample an ordered tuple $\mathbf{v}=(v_{i})_{i\in[k]}$ of distinct vertices in $[n]$ at random. Next, we define the drawn graph $F\in\mathcal{H}_{k}$ by $ij\in E(F)$ if and only if $v_{i}v_{j}\in E(G_{\ell-1})$ and generate the replacement graph $H\in\mathcal{H}_{k}$ according to the probability distribution $(\mathcal{R}_{F,H})_{H\in\mathcal{H}_{k}}$. Finally, we replace $G_{\ell-1}[\mathbf{v}]$ with $H$ to obtain $G_{\ell}$ from $G_{\ell-1}$.
Flip processes were studied in [7] through the lens of dense graph limits [9]. Let $(\Omega,\pi)$ be an atomless probability space with an implicit separable sigma-algebra. Write $\mathcal{W}\subseteq L^{\infty}(\Omega^{2})$ for the set of kernels, that is, bounded symmetric measurable real-valued functions. Write $\mathcal{W}_{0}$ for the set of graphons, that is, the elements of $\mathcal{W}$ whose range is a subset of $[0,1]$. An important metric on $\mathcal{W}$ is the cut norm distance $d_{\square}$ given by $d_{\square}(U,W)=\sup_{S,T\subseteq\Omega}\left|\int_{S\times T}(U-W)\,\mathrm{d}\pi^{2}\right|$. Let $G$ be a graph and $(\Omega_{i})_{i\in V(G)}$ be a partition of $\Omega$ into parts of equal measure. We write $W_{G}$ for the graphon which is equal to $\mathbb{I}_{\left\{ij\in E(G)\right\}}$ on $\Omega_{i}\times\Omega_{j}$ and call it a graphon representation of $G$.
@@
The distribution of the flip process started at a finite graph $G$ with rule $\mathcal{R}$ is its usual probability distribution from the theory of stochastic processes. We say that two rules $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ of orders $k_{1}$ and $k_{2}$ respectively have the same flip process distributions if for all finite graphs $G$ on $n\geq\max\{k_{1},k_{2}\}$ vertices the flip process started at $G$ with rule $\mathcal{R}_{1}$ has the same distribution as the flip process started at $G$ with rule $\mathcal{R}_{2}$. Note that two rules with the same trajectories could have different flip process distributions. A rule $\mathcal{R}$ is ignorant if the replacement graph distribution $(\mathcal{R}_{F,H})_{H\in\mathcal{H}_{k}}$ is independent of the drawn graph $F\in\mathcal{H}_{k}$. It was observed in [1, Section 4] that two ignorant rules $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ of the same order and with the same expected replacement edge count $d_{\mathcal{R}}=\sum_{H\in\mathcal{H}_{k}}\mathcal{R}_{F,H}e(H)$ have the same trajectories, that is, $\Phi_{\mathcal{R}_{1}}=\Phi_{\mathcal{R}_{2}}$. More concretely, consider ignorant rules $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ of order $4$ where for $\mathcal{R}_{1}$ the replacement graph is either complete or empty with probability $\frac{1}{2}$ each and for $\mathcal{R}_{2}$ the replacement graph is a copy of $K_{1,3}$. Clearly, they have different replacement behaviour. On the other hand, both rules are ignorant rules of the same order with the same expected replacement edge count $d_{\mathcal{R}_{1}}=d_{\mathcal{R}_{2}}=3$, so they have the same trajectories. Indeed, the lens of dense graph limits captures the macroscopic profile of flip processes but is insensitive to lower order fluctuations. This leads us to the following question.
@@
Here $T_{F^{a,b}}W(x,y)$ is invariant under the relabelling of vertices of $F^{a,b}$ (including $a$ and $b$), which yields an isomorphic rooted graph in the sense of the definitions on page 2. Hence, we may focus on equivalence classes of rooted graphs $F^{a,b}$ related by relabelling; write $\mathcal{J}_{k}$ for the set of such equivalence classes for $\mathcal{G}_{k}$. For each $J\in\mathcal{J}_{k}$ let $T_{J}W(x,y)$ be the unique value of $T_{F^{a,b}}W(x,y)$ for all $F^{a,b}\in J$ and set $q_{J,\mathcal{R}}=\sum_{F^{a,b}\in J}q_{F^{a,b},\mathcal{R}}$. Hence, we have
@@
Theorem 1.3 characterizes equivalence classes of rules of the same order with the same trajectories. Let us now demonstrate how it may be applied to determine whether two rules $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ of different orders have the same trajectories. Say the orders of $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are $k_{1}$ and $k_{2}$ respectively; by symmetry we may assume $k_{1}<k_{2}$. Let $\mathcal{R}^{*}_{1}$ be the rule of order $k_{2}$ where the replacement graph $H$ is obtained from the drawn graph $F$ by replacing the subgraph of $F$ induced on $k_{1}]$ according to rule $\mathcal{R}_{1}$ and leaving all other pairs unchanged. Since $\mathcal{R}^{*}_{1}$ acts exclusively on pairs on $k_{1}]$ and on those pairs it acts in accordance with $\mathcal{R}_{1}$, it follows that $\mathcal{R}^{*}_{1}$ and $\mathcal{R}_{1}$ have the same trajectories. Let $F^{a,b}$ be a rooted graph with $F\in\mathcal{H}_{k_{2}}$ and distinct $a,b\in[k_{2}]$. If $a,b\in[k_{1}]$, then we have $q_{F^{a,b},\mathcal{R}^{*}_{1}}=q_{G^{a,b},\mathcal{R}_{1}}$ with $G$ being the subgraph of $F$ induced on $k_{1}]$. Otherwise, we have $q_{F^{a,b},\mathcal{R}^{*}_{1}}=0$. Now we apply Theorem 1.3 to determine whether $\Phi_{\mathcal{R}^{*}_{1}}=\Phi_{\mathcal{R}_{2}}$, which is equivalent to $\Phi_{\mathcal{R}_{1}}=\Phi_{\mathcal{R}_{2}}$. All in all, this gives us a complete characterization of equivalence classes of flip process rules with the same trajectories.
Many combinatorially interesting rules turn out to be symmetric deterministic rules. Write $S_{k}$ for the symmetric group on $[k]$. We have the following natural action of $S_{k}$ on $\mathcal{H}_{k}$. Given $\sigma\in S_{k}$ and $F\in\mathcal{H}_{k}$ we define $\sigma\cdot F$ to be the graph with vertex set $[k]$ and edge set $\{\sigma(i)\sigma(j):ij\in E(F)\}$. A rule $\mathcal{R}$ of order $k$ is symmetric if for all $F,H\in\mathcal{H}_{k}$ and $\sigma\in S_{k}$ we have $\mathcal{R}_{\sigma\cdot F,\sigma\cdot H}=\mathcal{R}_{F,H}$ and deterministic if for each $F\in\mathcal{H}_{k}$ there is a unique $G\in\mathcal{H}_{k}$ such that $\mathcal{R}_{F,H}=\mathbb{I}_{\left\{H=G\right\}}$. Indeed, naturally arising rules are likely to act symmetrically on isomorphic graphs. For example, naturally arising rules of order $3$ are likely to treat the cherry graph $K_{1,2}$ symmetrically, no matter whether the middle vertex is labelled $1$, 22 or $3$. The class of symmetric deterministic rules contains many natural families of rules which were highlighted in [7] and studied in [1]. These include the complementing rules, the component completion rules, the extremist rules, and the clique removal rules.
@@ 2. Main Theorem
    There is an open interval $\mathcal{D}_{\mathcal{R},W}\subseteq\mathbb{R}$ containing 0 and a trajectory $\Phi_{\mathcal{R}}^{\cdot}W\colon\mathcal{D}_{\mathcal{R},W}\to\mathcal{W}$ starting at $W$ such that any other trajectory starting at $W$ is a restriction of $\Phi_{\mathcal{R}}^{\cdot}W$ to a subinterval of $\mathcal{D}_{\mathcal{R},W}$.
@@
    If $W\in\mathcal{W}_{0}$ is a graphon, then the set $L_{\mathcal{R},W}:=\{t\in\mathcal{D}_{\mathcal{R},W}:\Phi_{\mathcal{R}}^{t}W\in\mathcal{W}_{0}\}$ is a closed interval containing $0,\infty)$.
We introduce some further definitions. We say that rooted graphs $\mathbf{F}$ and $\mathbf{H}$ are isomorphic if there is an isomorphism from $\mathbf{F}$ to $\mathbf{H}$, that is, a bijective function $\phi\colon V(\mathbf{F})\to V(\mathbf{H})$ such that
  1. (F1)
    for all $u,v\in V(\mathbf{F})$ we have $uv\in E(\mathbf{F})$ if and only if $\phi(u)\phi(v)\in E(\mathbf{H})$,
  2. (F2)
    for all $v\in V(\mathbf{F})$ we have $\phi(v)\in R(\mathbf{H})$ if and only if $v\in R(\mathbf{F})$, and
  3. (F3)
    for all $u,v\in R(\mathbf{F})$ we have $\phi(u)\leq_{R(\mathbf{H})}\phi(v)$ if and only if $u\leq_{R(\mathbf{F})}v$;
write $\mathbf{F}\cong\mathbf{H}$. Let $k\in\mathbb{N}$. Write $\mathcal{J}_{k}$ for the set of isomorphism classes of the collection $\mathcal{G}_{k}$. For $J\in\mathcal{J}_{k}$ and a rule $\mathcal{R}$ of order $k$, set
@@
    For all $J\in\mathcal{J}_{k}$ we have $q_{J,\mathcal{R}_{1}}=q_{J,\mathcal{R}_{2}}$ .
@@
    $\mathfrak{V}_{\mathcal{R}_{1}}=\mathfrak{V}_{\mathcal{R}_{2}}$ on $\mathcal{W}_{0}$ .
@@
    $\mathfrak{V}_{\mathcal{R}_{1}}=\mathfrak{V}_{\mathcal{R}_{2}}$ on $\mathcal{W}$ .
@@
In this section we prove Theorem 2.3. We briefly describe our proof strategy here. Our proof scheme is (S1) $\Rightarrow$ (S3) $\Rightarrow$ (S4) $\Rightarrow$ (S2) $\Rightarrow$ (S1). We note that rooted induced densities are invariant under rooted graph isomorphism, so for $J\in\mathcal{J}_{k}$ we may set $T_{J}$ to be the unique operator equal to $T_{F^{a,b}}$ for all $F^{a,b}\in J$ and the first implication follows. The second implication follows from the definition and uniqueness of trajectories (Definition 2.1 and Theorem 2.2), while the third implication follows from the smoothness of the velocity operator (Lemmas 3.3 and 3.4). Indeed, the final implication (S2) $\Rightarrow$ (S1) is the most difficult step. The goal here is to show that the collection $\mathcal{I}=\{T_{J}:J\in\mathcal{J}_{k}\}$ of operators is linearly independent.
We apply a couple of key insights to achieve this goal. Firstly, we focus on parametrized graphon representations of graphs because evaluating induced densities at these graphons nicely discretizes them into sums of monomials indexed over functions between rooted graphs that satisfy (F1), (F2) and (F3). Secondly, we focus on twinfree versions of rooted graphs because Lemma 3.1 tells us that the aforementioned functions are necessarily injective when the graphs involved are twinfree; in particular, this induces a partial order on the set of twinfree versions of rooted graphs. Combining these two insights leads us to Lemma 3.2, which provides an explicit quantification of induced densities at parametrized graphon representations of graphs with a unique monomial assigned to each relevant rooted graph. Finally, the desired linear independence follows from the linear independence of monomials in polynomial functions (see Lemma 3.5).